120 is _______% of 40.
Answer by ptaylor(2198)

120 is _______% of 40.

Let's try an easier one first, ok?
2 is ________% of 4. In other words what number multiplied by 4 equals 2??
4*x=2; x=2/4 or 1/2
Well, we know right off that 2 is 1/2 of 4 and we also know that 1/2=0.5 and we also know that if we want to change a decimal to a percent we have to multiply by 100 so 0.5*100=50 or 50% and we knew that before we started, didn't we?
Now, lets look at your problem. What number multiplied by 40 gives 120?
Right off we know that it's 3, but we could find the answer using the equation:
40*x=120 divide both sides by 40
x=3
If we want to convert 3 to a percent, what do we do? We multiply by 100, so 3*100=300%!!!!!
120 is 300% of 40


Hope this helps---ptaylor